On the blueprint of Lauren's new office building, the scale is 3 in. equals 8 ft. What will be the actual length of Lauren's office if it measures 6 in. long on the blueprint?
_______________________________________________________________________________
Determine whether the pair of fractions is proportional: 2/5=6/15

Answer by stanbon(48526)

On the blueprint of Lauren's new office building, the scale is 3 in. equals 8 ft. What will be the actual length of Lauren's office if it measures 6 in. long on the blueprint?
---------------
Let x be the actual length of Lauren's office.
x/6 = 8/3
x = 6(8/3)
x = 2*8
x = 16 ft
=====================
Cheers,
Stan H.